Suppose a random sample of 38 sports cars has an average annual fuel cost of K2218 and the standard deviation was K523. Construct a 90% confidence interval for μ. Assume the annual fuel costs are normally distributed.


1
Expert's answer
2022-06-07T19:09:26-0400

Based on the information provided, the significance level is "\alpha = 0.10,"  and the critical value for a two-tailed test is "z_c =1.6449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}})""=(2218-1.6449\times\dfrac{523}{\sqrt{38}},2218+1.6449\times\dfrac{523}{\sqrt{32}})""=(2078.4437, 2357.5563)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "2078.4437 < \mu < 2357.5563," which indicates that we are 90% confident that the true population mean "\mu"  is contained by the interval "(2078.4437, 2357.5563)."